OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1

OpenStudy (anonymous):

First solve for y in the first equation. 2x + y = 3 y = -2x + 3 Then substitute this into the second equation for y. -3x + 4y = 1 -3x + 4(-2x + 3) = 1 -3x - 8x + 12 = 1 After you solve for x, plug it back in to solve for y. Can you solve from here?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (anonymous):

You're welcome :)